# Dependence of the critical temperature and disorder in holographic superconductors on superfluid density

### v3

@@ Abstract
Recent experiments strongly indicate deep connections between transports of strange metal and high $T_{c}$ superconductors. For example, the dependence of the zero-temperature phase stiffness on the critical superconducting temperature is generally linear, which is incompatible with the standard Bardeen-Cooper-Schrieffer description. We develop an analytical method for AC conductivity calculation and explore the scaling relations among superconducting critical temperature, superfluid density, and momentum dissipation strength for the Gubser-Rocha model with extensions in the probe limit. In the normal phase, we show that the critical temperature is proportional to the momentum dissipation strength in a certain parameter range, which is universal in holographic models. In the superconducting phase, studying the AC conductivity analytically and numerically, we find linear dependence of zero-temperature superfluid density (phase stiffness) on the critical superconducting temperature, which is consistent with recent experiments of high $T_{c}$ superconductors. These results further underpin the deep connections between strange metal and high $T_{c}$ superconductors.
@@ 1 Introduction
Superconductivity is a very significant discovery in physics in the early part of the 20th century. Since its ascertainment, overwhelming experimental and theoretical research has been done on superconductivity, one of the most striking is BCS theory[1]. This theory assumes that the fundamental characteristics of superconductivity are caused by Cooper’s pair correlation. As experiments progressed, high-temperature oxide superconductors[2, 3, 4, 5] and iron-based superconductors[6, 7, 8, 9] were found. The description of these superconductors is beyond the scope of BCS theory. Recently, a series of groundbreaking experimental findings have emerged that demonstrate how the superfluid density increasingly manifests a linear relationship with temperature as temperatures escalate [10, 11, 12]. This phenomenon departs from the conventional expectations set by the Bardeen-Cooper-Schrieffer (BCS) theory, thereby shedding light on potentially novel facets of high-temperature superconductivity and its fundamental physics. There are a lot of interests and challenges in studying the critical temperature $T_{c}$ related to the superfluid density $n_{s}$. Some experimental papers [13, 15, 14] elucidate that $T_{c}$ seems to be principally controlled by the superfluid density. More remarkably, the superfluid density displays a strong linear temperature dependence within a certain temperature range.
@@
In this paper, we explore the universal properties of the superconductors, e.g., scaling relations, by using the AdS/CFT correspondence in a specific holographic model. The AdS/CFT correspondence[16, 18, 17], also known as holography or gauge/gravity duality, is a framework for tracing the behavior of strongly-coupled quantum systems in terms of a higher dimensional gravity theory. It can be a very powerful and fascinating tool to tackle strongly interacting systems by studying gravitational models. In particular, holographic techniques have been widely applied for studying condensed matter physics [19, 21, 20, 22].
@@
The kind of model called holographic superconductors is still an attractive topic of theoretical research. They have the potential to provide new insights into the nature of superconducting materials and could lead to the development of new technologies based on superconductivity. Based on the AdS/CFT correspondence, the first minimal bottom-up construction had been implemented by Hartnoll, Herzog, and Horowitz[23]. In Ref. [23], the authors considered the charged complex scalar and the Maxwell field in the Schwarzschild-AdS4 (SAdS4) spacetime. Although this is a simple model, it can reproduce the typical behavior of the superconductor successfully. The calculation done in [23] was the so-called probe limit, ignoring the backreactions to the gravity sector from the matter sector. In [24], the authors considered the Einstein-Maxwell theory with the complex scalar and found that the probe-limit calculation corresponds to the infinite charge case with appropriate scaling. In any case, the complete analysis of the superconducting phase requires numerical computations. Meanwhile, some analytic properties of the holographic superconductor in the superconducting phase have been studied[26, 25, 27]. Since the complete analysis requires numerical computations, they provided some approximate formulas. From then on [23, 24], many papers have appeared that attempt to apply this useful idea to investigate superconductors in different setups [28, 31, 29, 30, 34, 35, 32, 33, 36].
@@
Recently, it has been recognized that the strange metallic behavior is universal in superconductors like the high-$T_{c}$ cuprates. The strange metal is often described by its linear-$T$ dependence of the resistivity[37, 38, 39]. In holographic studies, the Gubser-Rocha model[40] is known as one of the candidates describing the strange metallic behavior of the superconductors. This model is an Einstein-Maxwell-dilaton theory, and axion fields are often introduced to break the translational invariance[41, 42]. The original model is consistent with the specific truncation of the 11-dimensional supergravity. This model has been widely utilized for studying the holographic version of the high-$T_{c}$ cuprates superconductors. On the other hand, we notably advance the analytical framework for studying holographic superconductors by developing novel approximate methods to explore key analytic properties within the context of the Gubser-Rocha model. Despite the prevalent reliance on numerical computations in previous studies, akin to those conducted in the foundational holographic superconductor paradigm, we focus our efforts here on investigating analytically significant aspects such as the critical temperature and alternating current (AC) conductivity.
@@
The primary objective of this paper is to investigate the scaling relationships connecting the superconducting critical temperature, the superfluid density, and the momentum-dissipation strength within the framework of the holographic superconductor model derived from the Gubser-Rocha model. Moreover, a significant part of our study involves comparing these theoretically deduced relations with empirical results obtained in experimental settings. First, we study the critical temperature by finding the onset of the charged-scalar instability. Taking advantage of the Strum-Liouville method, we investigate how the momentum-dissipation strength affects the critical temperature. The results obtained by numerical and approximate methods are in good agreement with each other. Using the approximate formula, we find the scaling relation between the critical temperature and the strength of the momentum dissipation. Subsequently, we delve into the study of the AC conductivity in the superconducting state. To extract the superfluid density, we meticulously analyze the AC conductivity data and proceed to examine the correlation between the zero-temperature superfluid density and the critical temperature. Consequently, our findings reveal a linear scaling relationship between these parameters in a particular scenario, which strikingly aligns with recent experimental outcomes observed in copper oxides [15]. Finally, we develop and scrutinize an analytical approximate formula for the AC conductivity specifically tailored for the superconducting phase.
In brief, this paper is driven by three core motivations and objectives:
Unveiling Universal Behavior: The first pivotal objective centers around the exploration of the universal characteristics of the superfluid density within our holographic superconductor model. The main findings are encapsulated in Fig. 8 and Eq. (48), which represent the cornerstone results of this study. To substantiate this relationship across the entire temperature range within the superconducting phase, it necessitates the meticulous computation of several fundamental superconductor properties within our theoretical framework.
Implementing Probe-Limit Calculations: Throughout the paper, we effectively achieve the secondary aim through demonstrating the probe-limit calculation within our holographic model grounded on the Gubser-Rocha model. This endeavor not only serves as a methodological achievement but also holds significant physical implications for developing more realistic models of high-temperature superconductors. As an example, the extension to d-wave superconductors can be readily explored by substituting the appropriate matter sector. Thus, our research constitutes a foundational step towards considering such extensions.
Analyzing Analytic Results: For the third goal, equations (32) and (59) correspondingly highlight the analytic outcomes of our study. Notably, the linear relation depicted in equation (32) appears to exhibit universality across various holographic models incorporating momentum-dissipation strength. Despite securing only a limited number of analytical results, this investigation underscores the challenges and limitations inherent to analytical studies within our specific model, thereby contributing to a deeper understanding of its complexity and scope.
@@
The organization of this paper is as follows. The holographic setup of our model is given in section 2. We investigate the critical temperature of the superconducting phase in section 3. In section 4, we turn to study the conductivity, and consider an analytic approximation for the AC conductivity in the superconducting phase. We provide conclusion and discussion in section 5. In Appendix A, we demonstrate how the profile of the critical temperature with general $q$ from the full analysis with backreactions agrees with the probe limit results in a large $q$ limit. In Appendix B, we show another approximate formula for the AC conductivity.
@@ 2 Setup
In this study, we employ the Gubser-Rocha model [40] extended with the linear axions and the charged scalar, for studying the superconducting nature [43, 44, 45, 46]. The action of our model is given by [45, 46]
@@
The action consists of two components, $\mathcal{L}_{g}$ representing the Lagrangian of fields involving the gravity and $\mathcal{L}_{m}$ representing the Lagrangian of matter fields. The model involves gravity fields $g_{\mu\nu}$, a dilaton $\phi$, axions $\psi_{I}$, $U(1)$ gauge fields and a complex scalar field $\Phi$. It was shown that the presence of the dilaton field can realize the vanishing entropy at zero temperature[47]. The axions are originally introduced to break the translational invariance explicitly in the full analysis with backreactions [43]. $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength of the $U(1)$ gauge field. The covariant derivative $D$ is defined by $D_{\mu}:=\nabla_{\mu}-iqA_{\mu}$. $B(\phi)$ is a coupling factor between dilaton field $\phi$ and complex scalar field $\Phi$.
We concentrate our analysis on the probe limit, which corresponds to taking the limit as q approaches infinity. Under this particular condition, we can confidently assert that the gravity $\mathcal{L}_{g}$ and matter sectors $\mathcal{L}_{m}$ are effectively decoupled and described independently from one another. The equations of motion for the dilaton and the axions are obtained from $\mathcal{L}_{g}$ as
@@
we obtain the following solution [43]
@@
where $P$ is an integration constant corresponding to the location of the curvature singularity at $r=P$ and it can be expressed in terms of temperature $T$ and momentum dissipation strength $k$, see Eq. (6).111 Note that the blackening factor is still different from $f(r)=r^{2}(1-r_{h}^{3}/r^{3})$ in the SAdS4, even if one sets $P=0$. $k$ is a positive constant of the linear axions, which can be understood as the strength of the momentum dissipation in the backreacted setup. Remark that $k$ was introduced to break the translational symmetry. In this study, we naively expect that $k$ still holds some aspects of the strength of the momentum dissipation, even in the probe limit. The geometry is a black hole spacetime with the horizon at $r=r_{h}=k/\sqrt{2}$. To avoid the naked singularity, the range of $P$ is restricted by $-\infty<P<r_{h}$. In the probe limit, the gravity sector is decoupled from the Maxwell field. Therefore, this solution is a neutral black hole solution. Note that this solution was also studied in Refs. [49, 48], and Ref. [50] without the axions. This is also a special neutral case of the charged black hole studied in Refs. [43, 51, 44, 45, 46]. (See also Appendix A). The authors of Ref. [46] studied the same model for general $q$ beyond the probe limit. We call the above solution as ‘dilatonic black hole’. The Hawking temperature reads
$T=\frac{r_{h}}{2\pi}\sqrt{1-\frac{P}{r_{h}}}.$ (6)
In order to obtain positive $T$, the range of $P$ is restricted in $-\infty<P<r_{h}$.
On the other hand, the model also admits another neutral solution without the dilaton hair, which is given by [41]
@@
We call it as ‘bald black hole’ solution in this study. In this solution, $k$ and $r_{h}$ are independent parameters but the number of parameters does not change. Since the dilaton vanishes, the solution is same as the neutral solution in the linear axions model [41]. Now, the Hawking temperature is given by
$T=\frac{1}{8\pi}\frac{6r_{h}^{2}-k^{2}}{r_{h}}.$ (8)
The range of $k$ is restricted by $T>0$ as $k<\sqrt{6}r_{h}$. This solution is thermodynamically favored in the range $T>\frac{k}{2\sqrt{2}\pi}$ [44, 48]. Therefore, we need to consider this solution as the background metric rather than the dilatonic black hole (5), at high temperatures where $T>k/(2\sqrt{2}\pi)$.
@@
where $M$ is a mass of the scalar field near $r=\infty$, and $\tau$ is a constant determining the coupling with the dilaton field. In this study, we fix $M^{2}=-2$, which leads the dimension of the charged-scalar operator as $\Delta=1$ or $2$. The authors of Ref. [46] indicate that the superconducting instability can be triggered easily at higher coupling $\tau$. For more detail argument about $B(\phi)$, see Ref. [52]. Now, we consider the following ansatz:
@@
For the vector field, $\mu$ and $\rho$ are dual to the chemical potential and charge density of the boundary conformal field theory, respectively. For the scalar field, both the leading and the subleading terms can be normalizable in our case [53]. Similarly to Ref. [23], we refer the case when $\Phi^{(i)}=\expectationvalue{O_{i}}$ and $\Phi^{(j)}=0$ as $O_{i}$ theory, where $i=1$ or $2$ and $j\neq i$. The choice of the $O_{2}$ theory is called standard quantization, whereas the $O_{1}$ theory is realized by adding corresponding boundary action. $\expectationvalue{O_{i}}$ is considered as an operator with dimension $i$ in the boundary theory.
@@
3 Scaling relation between $T_{c}$ and $k$

### v4

# 2312.16029

@@ Abstract
Recent experiments strongly indicate deep connections between transports of strange metal and high $T_{c}$ superconductors. For instance, it is known that the dependence of the zero-temperature phase stiffness on the critical superconducting temperature becomes linear in underdoped materials. In this paper, we investigate relation meticulously between the phase stiffness and the critical superconducting temperature for the Gubser-Rocha holographic superconductor model in the probe limit. The superfluid density (or phase stiffness) can be extracted from the low-frequency dependence of the AC conductivity. More importantly, we find linear dependence of zero-temperature superfluid density on the critical superconducting temperature, which has a similarity to a recent experiment in a film cuprate. In addition, we also provide some approximate formulas for the critical temperatures and the AC conductivity.
@@ 1 Introduction
Superconductivity is a very significant discovery in physics in the early part of the 20th century. Since its ascertainment, overwhelming experimental and theoretical research has been done on superconductivity, one of the most striking is Bardeen-Cooper-Schrieffer (BCS) theorybcs . This theory assumes that the fundamental characteristics of superconductivity are caused by Cooper’s pair correlation. As experiments progressed, high-temperature oxide superconductorsmuller ; mkwu ; hiroshi ; zzsheng and iron-based superconductorsjjhamlin ; kamihara ; xhchen ; haihuwen were found. The description of these superconductors is beyond the scope of BCS theory. Recently, a series of groundbreaking experimental findings have emerged that demonstrate how the superfluid density increasingly manifests a linear relationship with temperature as temperatures escalate mm ; soumyajit ; sudhansu . These phenomena depart from the conventional expectations set by the BCS theory, thereby shedding light on potentially novel facts of high-temperature superconductivity and its fundamental physics. There are lots of interests and challenges in studying the critical temperature $T_{c}$ related to the superfluid density $n_{s}$. Some experimental papers lulian ; copperoxides ; microwavetl2201 elucidate that $T_{c}$ seems to be principally controlled by the superfluid density. More remarkably, the superfluid density displays a strong linear temperature dependence within a certain temperature range. In Ref. copperoxides , the authors have been reported the phase stiffness depend on temperature and doping by investigating the entire overdoped side of the $\text{La}_{2-x}\text{Sr}_{x}\text{CuO}_{4}$ (LSCO) phase diagram. In their results, the dependence of the zero-temperature phase stiffness on the critical is mostly linear, but with an offset. However, close to the origin this dependence becomes parabolic. This phenomenon was taken as evidence for the violation of the Homes’ law. In Ref. dordevic , the authors have unraveled explicitly that the Homes’ law is indeed violated, and as the main reason for violation is that the superfluid density in LSCO films is suppressed more strongly than in bulk samples.
@@
Based on the above experimental findings, in this paper, we explore the universal properties of the superconductors, e.g., scaling relations, specially concentrate on the dependence of the zero-temperature phase stiffness on the critical superconducting temperature, by using the AdS/CFT correspondence in a specific holographic model. The AdS/CFT correspondenceAdSCFT ; Gubser:1998bc ; Witten:1998qj , also known as holography or gauge/gravity duality, is a framework for tracing the behavior of strongly-coupled quantum systems in terms of a higher dimensional gravity theory. It can be a very powerful and fascinating tool to tackle strongly interacting systems by studying gravitational models. In particular, holographic techniques have been widely applied for studying condensed matter physics zaanen ; Baggioli:2019rrs ; ammon ; Cai:2015cya .
@@
The kind of model called holographic superconductors is still an attractive topic of theoretical research. They have the potential to provide new insights into the nature of superconducting materials and could give rise to the development of new technologies based on superconductivity. Based on the AdS/CFT correspondence, the first minimal bottom-up construction had been implemented by Hartnoll, Herzog, and Horowitzbhs . In Ref. bhs , the authors considered the charged complex scalar and the Maxwell field in the Schwarzschild-AdS4 (SAdS4) spacetime. Although this is a simple model, it can reproduce the typical behavior of the superconductor successfully. The calculation done in bhs was the so-called probe limit, ignoring the backreactions to the gravity sector from the matter sector. In Hartnoll:2008kx , the authors considered the Einstein-Maxwell theory with the complex scalar and found that the probe-limit calculation corresponds to the infinite charge case with appropriate scaling. In any case, the complete analysis of the superconducting phase requires numerical computations. Meanwhile, some analytic properties of the holographic superconductor in the superconducting phase have been studiedHerzog:2009xv ; Gary ; Siopsis:2010uq ; Ge:2010aa . Since the complete analysis requires numerical computations, they provided some approximate formulas. From then on bhs ; Hartnoll:2008kx , many papers have appeared that attempt to apply this useful idea to investigate superconductors in different setups Franco:2009yz ; Gangopadhyay:2013qza ; Cai:2010cv ; Roychowdhury:2012hp ; Sheykhi:2016kqh ; Qiao:2020hkx ; Erdmenger:2013zaa ; Ge:2015fmu ; Zhao:2023qms .
@@
Recently, it has been recognized that the strange metallic behavior is universal in superconductors like the high-$T_{c}$ cuprates. The strange metal is often described by the linear-$T$ dependence of its resistivityyuan ; jiang ; Phillips:2022nxs . In holographic studies, the Gubser-Rocha modelGubser:2009qt is known as one of the candidates describing the strange metallic behavior of the superconductors. This model is an Einstein-Maxwell-dilaton theory, and axion fields are often introduced to break the translational invarianceAndrade:2013gsa ; Gouteraux:2014hca . The original model is consistent with the specific truncation of the 11-dimensional supergravity. We can explore microscopic mechanism within this model such as quantum critical points and superconducting transition temperature, which help us understand the origin of high-temperature superconductivity more profoundly. This model has been widely utilized for studying the holographic version of the high-$T_{c}$ cuprates superconductors. As well known, the Homes’ law is a universal relation of superconductors between the superfluid density $n_{s}$ at zero temperature, the cirtical temperature $T_{c}$ and the direct current (DC) conductivity $\sigma_{DC}$, which is observed in high $T_{c}$ superconductors. In Ref. Homes' law , the authors have investigated how Home’s law can be realized together with linear-$T$ resistivity by using holographic models related to the Gubser-Rocha model. Inspired by literature copperoxides ; dordevic , our main goal is to study the phenomena that deviate from Holmes’ law. Because there has been no study for such phenomenon with alternating current (AC) conductivity in holography, our study is a important and essential step to look into the novel experimental results. On the other hand, we notably advance the analytical framework for studying holographic superconductors by developing novel approximate methods to explore key analytic properties within the context of the Gubser-Rocha model. Despite the prevalent reliance on numerical computations in previous studies, akin to those conducted in the foundational holographic superconductor paradigm, we focus our efforts here on investigating analytically significant aspects such as the critical temperature and AC conductivity.
@@
The primary objective of this paper is to investigate the scaling relationships connecting the superconducting critical temperature, the superfluid density, and the momentum-dissipation strength within the framework of the holographic superconductor model derived from the Gubser-Rocha model. Moreover, a significant part of our study involves comparing these theoretically deduced relations with empirical results obtained in experimental settings. First, we study the critical temperature by finding the onset of the charged-scalar instability. Taking advantage of the Sturm-Liouville method, we investigate how the momentum-dissipation strength affects the critical temperature. The results obtained by numerical and approximate methods are in good agreement with each other. Using the approximate formula, we find the scaling relation between the critical temperature and the strength of the momentum dissipation. Subsequently, we compute the AC conductivity in the superconducting state. To extract the superfluid density, we meticulously analyze the AC conductivity data and proceed to examine the correlation between the zero-temperature superfluid density and the critical temperature. Most notably, our results roughly show a linear relationship between the zero-temperature phase stiffness and the critical temperature. This result is similar to recent experimental outcomes observed in a certain overdoped copper-oxide film copperoxides . Interestingly, this material does not obey the Homes’ law dordevic .
@@
The organization of this paper is as follows. The holographic setup of our model is given in section 2. We investigate the critical temperature of the superconducting phase in section 3. In section 4, we turn to study the AC conductivity for obtaining the superfluid density. We provide conclusion and discussion in section 5. In Appendix A, we demonstrate how the profile of the critical temperature with general $q$ from the full analysis with backreactions agrees with the probe limit results in a large $q$ limit. In Appendix B, we show some results of the approximate formula for the AC conductivity in the ordered phase.
@@ 2 Setup
In this study, we employ the Gubser-Rocha model Gubser:2009qt extended with the linear axions and the charged scalar, for studying the superconducting nature Zhou:2015qui ; Kim:2017dgz ; Jeong:2018tua ; Homes' law . Particularly, we study the relation between superfluid density $n_{s}$ and superconducting critical temperature $T_{c}$ which has not been explored so far. The action of our model is given by Jeong:2018tua ; Homes' law
@@
The action consists of two components, $\mathcal{L}_{g}$ representing the Lagrangian of fields involving the gravity and $\mathcal{L}_{m}$ representing the Lagrangian of matter fields. The model involves gravity fields $g_{\mu\nu}$, a dilaton $\phi$, axions $\psi_{I}$, $U(1)$ gauge fields and a complex scalar field $\Phi$. It was shown that the presence of the dilaton field can realize the vanishing entropy at zero temperatureDavison:2013txa . The axions are originally introduced to break the translational invariance explicitly in the full analysis with backreactions Zhou:2015qui . $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength of the $U(1)$ gauge field. The covariant derivative is defined by $D_{\mu}:=\nabla_{\mu}-iqA_{\mu}$. $B(\phi)$ is a coupling factor between dilaton field $\phi$ and complex scalar field $\Phi$.
We concentrate our analysis in the probe limit, which corresponds to taking the limit as q approaches infinity. Under this particular condition, we can confidently assert that the gravity $\mathcal{L}_{g}$ and matter sectors $\mathcal{L}_{m}$ are effectively decoupled and described independently from one another. The equations of motion for the dilaton and the axions are obtained from $\mathcal{L}_{g}$ as
@@
we obtain the following solution Zhou:2015qui
@@
where $P$ is an integration constant corresponding to the location of the curvature singularity at $r=P$ and it can be expressed in terms of temperature $T$ and momentum dissipation strength $k$, see Eq. (6).111 Note that the blackening factor is still different from $f(r)=r^{2}(1-r_{h}^{3}/r^{3})$ in the SAdS4, even if one sets $P=0$. $k$ is a positive constant of the linear axions, which can be understood as the strength of the momentum dissipation in the backreacted setup. Remark that $k$ was introduced to break the translational symmetry. In this study, we naively expect that $k$ still holds some aspects of the strength of the momentum dissipation, even in the probe limit. The geometry is a black hole spacetime with the horizon at $r=r_{h}=k/\sqrt{2}$. To avoid the naked singularity, the range of $P$ is restricted by $-\infty<P<r_{h}$. In the probe limit, the gravity sector is decoupled from the Maxwell field. Therefore, this solution is a neutral black hole solution. Note that this solution was also studied in Refs. Ren:2021rhx ; Jeong:2023ynk , and Ref. Ren:2022qkr without the axions. This is also a special neutral case of the charged black hole studied in Refs. Zhou:2015qui ; Caldarelli:2016nni ; Kim:2017dgz ; Jeong:2018tua ; Homes' law . (See also Appendix A). The authors of Ref. Homes' law studied the same model for general $q$ beyond the probe limit. We call the above solution as ‘dilatonic black hole’. The Hawking temperature and the entropy density read
$T=\frac{r_{h}}{2\pi}\sqrt{1-\frac{P}{r_{h}}},\quad s=4\pi r_{h}^{2}\sqrt{1-\frac{P}{r_{h}}},$ (6)
respectively. In order to obtain positive $T$, the range of $P$ is restricted in $-\infty<P<r_{h}$. We can write the entropy density $s=4\sqrt{2}\pi^{2}kT$.
On the other hand, the model also admits another neutral solution without the dilaton hair, which is given by Andrade:2013gsa
@@
We call it as ‘bald black hole’ solution in this study. In this solution, $k$ and $r_{h}$ are independent parameters but the number of parameters does not change. Since the dilaton vanishes, the solution is same as the neutral solution in the linear axions model Andrade:2013gsa . Now, the Hawking temperature and the entropy density read
$T=\frac{1}{8\pi}\frac{6r_{h}^{2}-k^{2}}{r_{h}},\quad s=4\pi r_{h}^{2},$ (8)
respectively. The range of $k$ is restricted as $k<\sqrt{6}r_{h}$ by $T>0$. This solution is favored in the range $T>\frac{k}{2\sqrt{2}\pi}$ thermodynamicallyKim:2017dgz ; Jeong:2023ynk . Therefore, we need to consider this solution as the background metric rather than the dilatonic black hole (5), at higher temperature where $T>k/(2\sqrt{2}\pi)$.
@@
where $M$ is a mass of the scalar field near $r=\infty$, and $\tau$ is a constant determining the coupling with the dilaton field. In this study, we fix $M^{2}=-2$, which leads the dimension of the charged-scalar operator as $\Delta=1$ or $2$. The authors of Ref. Homes' law indicate that the superconducting instability can be triggered easily at higher coupling $\tau$. For more detail argument about $B(\phi)$, see Ref. Cremonini:2016bqw . Now, we consider the following ansatz:
@@
For the vector field, $\mu$ and $\rho$ are dual to the chemical potential and charge density of the boundary conformal field theory, respectively. For the scalar field, both the leading and the subleading terms can be normalizable in our case Witten:2001ua . Similarly to Ref. bhs , we refer the case when $\Phi^{(i)}=\expectationvalue{O_{i}}$ and $\Phi^{(j)}=0$ as $O_{i}$ theory, where $i=1$ or $2$ and $j\neq i$. The choice of the $O_{2}$ theory is called standard quantization, whereas the $O_{1}$ theory is realized by adding corresponding boundary action. $\expectationvalue{O_{i}}$ is considered as an operator with dimension $i$ in the boundary theory.
@@
3 Critical temperature and condensation